### 29 Dec 2023 v1

# AIJack: Security and Privacy Risk Simulator for Machine Learning

Abstract
This paper introduces AIJack, an open-source library designed to assess security and privacy risks associated with the training and deployment of machine learning models. Amid the growing interest in big data and AI, advancements in machine learning research and business are accelerating. However, recent studies reveal potential threats, such as the theft of training data and the manipulation of models by malicious attackers. Therefore, a comprehensive understanding of machine learning’s security and privacy vulnerabilities is crucial for the safe integration of machine learning into real-world products. AIJack aims to address this need by providing a library with various attack and defense methods through a unified API. The library is publicly available on GitHub (https://github.com/Koukyosyumei/AIJack).
1 Introduction
Machine learning (ML) has become a foundational component of diverse applications, spanning image recognition to natural language processing. As these technologies proliferate, the need to comprehend and address security risks associated with ML models becomes imperative.
While ML models continuously enhance accuracy, attackers can significantly diminish it by introducing artificially created data. Evasion Attacks [1, 2], using specialized algorithms, can induce malfunctions in models. Consider a road sign classification model; by adding imperceptible noise to a ”speed limit 30km” sign, it may misclassify it as ”speed limit 60km.” Another attack, Poisoning Attack [3], injects contaminated data during training, lowering model accuracy.
To counter such attacks, various strategies have been proposed. Certified robustness [4, 5], a technique to formally guarantee that adversarial examples cannot lead to undesirable predictions, has gained attention. Another approach is debugging machine learning models [6, 7, 8], which aims to identify inputs causing unexpected behaviors.
ML also poses privacy challenges. Collecting large amounts of data for training can infringe on privacy, leading to data breaches. Model Inversion Attacks [9, 10] reconstruct training data from pre-trained models, threatening sensitive information. Similar attacks, like Membership Inference Attacks [11], aim to determine if a data point is part of the model’s training data.
Privacy protection techniques, including differential privacy [12], k-anonymity [13], homomorphic encryption [14], and distributed learning [15, 16], have been proposed. Differential privacy prevents individual data inference, while homomorphic encryption enables arithmetic operations on encrypted data. Distributed methods like Federated Learning (FL) [15] and Split Learning [16] facilitate collaborative learning among data owners.
It is crucial to assess the security and privacy risks of ML models and evaluate countermeasure effectiveness. To simplify such simulations, we propose the open-source software, AIJack, offering various attack and defense methods. AIJack enables experimenting with various combinations of attacks and defenses with simple code. Built on PyTorch [17] and scikit-learn [18], users can easily incorporate AIJack into existing code.
2 Package Design
Type Subcategory Method
Collaborative Horizontal FL FedAVG [15], FedProx [19], FedMD [20], FedGEMS [21], DSFL [22], MOON [23], FedEXP [24]
Collaborative Vertical FL SplitNN [16], SecureBoost [25]
Attack Model Inversion MI-FACE [9], DLG [10], iDLG [26], GS [27], CPL [28], GradInversion [29], GAN Attack [30]
Attack Label Leakage Norm Attack [31]
Attack Poisoning History Attack [32], Label Flip [32], MAPF [32], SVM Poisoning [3]
Attack Backdoor DBA [33], Model Replacement [34]
Attack Free-Rider Delta-Weight [35]
Attack Evasion Gradient-Descent Attack [1], FGSM [2], DIVA [36]
Attack Membership Inference Shadow Attack [11]
Defense Homomorphic Encryption Paillier [14]
Defense Differential Privacy DPSGD [12], AdaDPS [37], DPlis [38]
Defense Anonymization Mondrian [13]
Defense Robust Training PixelDP [4], Cost-Aware Robust Tree Ensemble [5]
Defense Debugging Model Assertions [6], Rain [7], Neuron Coverage [8]
Defense Others Soteria [39], FoolsGold [40], MID [41], Sparse Gradient [42]
Table 1: List of Supported Algorithms
AIJack is designed with the following principles:
  • All-around abilities for both attack and defense: AIJack provides a flexible API for over 40 attack and defense algorithms (see Tab. 1 for the comprehensive lists). Users can experiment with various combinations of these methods.
  • PyTorch-friendly design: AIJack supports many PyTorch models, allowing integration with minimal modifications to the original code.
  • Compatibility with scikit-learn: AIJack supports many scikit-learn models, enabling easy integration of attacks and defenses.
  • Fast Implementation with C++ backend: AIJack employs a C++ backend for components like Differential Privacy and Homomorphic Encryption, enhancing scalability.
  • MPI-Backend for Federated Learning: AIJack supports MPI-backed Federated Learning for deployment in High-Performance Computing systems.
  • Extensible modular APIs: AIJack comprises simple modular APIs, allowing easy extension with minimal effort.
Code 1 and Code 2 show the example codes implementing Evasion Attack and Poisoning Attack against SVM trained with scikit-learn, respectively.
1from aijack.attack import (
2    Evasion_attack_sklearn
3)
4
5attacker = Evasion_attack_sklearn(
6    clf, initial_data_positive
7)
8malicious_data, log = attacker.attack(
9    initial_data_negative
10)
Code 1: Evasion Attack against SVM
1from aijack.attack import (
2    Poison_attack_sklearn
3)
4
5attacker = Poison_attack_sklearn(
6    clf, X_train, y_train
7)
8malicious_data, log = attacker.attack(
9    initial_data, 1, X_valid, y_valid
10)
Code 2: Poisoning Attack against SVM
Example codes for Federated Learning can be found in Code 3 to Code 6. Code 3 implements FedAVG with PyTorch, the standard Federated Learning protocol, Code 4 implements a Gradient Inversion Attack against Federated Learning (a type of Model Inversion Attack), Code 5 applies Paillier Encryption, one of the most popular Homomorphic Encryption methods, to mitigate inversion attacks, and Code 6 migrates Code 3 to MPI-backend.
1from aijack.collaborative import (
2  FedAVGClient,
3  FedAVGServer,
4  FedAVGAPI
5)
6
7
8
9
10
11
12
13
14
15clients = [FedAVGClient(model_1),
16           FedAVGClient(model_2)]
17server = FedAVGServer(clients, model_3)
18
19api = FedAVGAPI(
20    server, clients,
21    criterion, optimizers, dataloaders,
22)
23api.run()
Code 3: Standard Federated Learning (FL)
1from aijack.collaborative import (
2  FedAVGClient,
3  FedAVGServer,
4  FedAVGAPI
5)
6from aijack.attack import (
7  GradientInversionAttackServerManager
8)
9
10mg = GradientInversionAttackServerManager(
11    input_shape,
12)
13MaliciousServer = mg.attach(FedAVGServer)
14
15clients = [FedAVGClient(model_1),
16           FedAVGClient(model_2)]
17server = MaliciousServer(clients, model_3)
18
19api = FedAVGAPI(
20    server, clients,
21    criterion, optimizers, dataloaders,
22)
23api.run()
Code 4: Gradient Inversion Attack against FL
1from aijack.collaborative import (
2  FedAVGClient,
3  FedAVGServer,
4  FedAVGAPI
5)
6from aijack.defense import (
7  PaillierGradientClientManager,
8  PaillierKeyGenerator
9)
10
11keygenerator = PaillierKeyGenerator(512)
12pk, sk = keygenerator.generate_keypair()
13
14mg = PaillierGradientClientManager(pk, sk)
15PaillierClient = mg.attach(FedAVGClient)
16
17clients = [PaillierClient(model_1),
18           PaillierClient(model_2)]
19server = FedAVGServer(clients, model_3)
20
21api = FedAVGAPI(
22    server,
23    clients,
24    criterion,
25    optimizers,
26    dataloaders,
27)
28api.run()
Code 5: FL with Paillier Encryption
1from aijack.collaborative import (
2  FedAVGClient,
3  FedAVGServer,
4  MPIFedAVGAPI,
5  MPIFedAVGClientManager,
6  MPIFedAVGServerManager
7)
8
9mcm = MPIFedAVGClientManager()
10msm = MPIFedAVGServerManager()
11MPIClient = mcm.attach(FedAVGClient)
12MPIGServer = msm.attach(FedAVGServer)
13
14if myid == 0:
15    client_ids = list(range(1, size))
16    server = MPIServer(comm, client_ids, model)
17    api = MPIFedAVGAPI(
18        comm, server, True, criterion
19    )
20else:
21    client = MPIClient(comm, model, user_id=myid)
22    api = MPIFedAVGAPI(
23        comm, client, False, criterion,
24        optimizer, dataloader
25    )
26
27api.run()
Code 6: FL with MPI backend
Code7 illustrates the standard workload involved in training a neural network using Differential Private Stochastic Gradient Descent (DPSGD). Our API’s design draws inspiration from Opacus [43], a renowned framework dedicated to deep learning with differential privacy. Additionally, our AIJack facilitates a more intuitive implementation, closely aligning with the algorithm initially proposed in[12].
1from aijack.defense import GeneralMomentAccountant, DPSGDManager
2
3accountant = GeneralMomentAccountant(noise_type="Gaussian", backend="cpp")
4privacy_manager = DPSGDManager(accountant, optim.SGD)
5
6accountant.reset_step_info()
7dpoptimizer_cls, lot_loader, batch_loader = privacy_manager.privatize(
8    noise_multiplier=sigma
9)
10optimizer = dpoptimizer_cls(net.parameters(), lr=lr)
11
12for epoch in range(num_iterations):
13    for X_lot, y_lot in lot_loader(optimizer):
14        for X_batch, y_batch in batch_loader(TensorDataset(X_lot, y_lot)):
15            optimizer.zero_grad()
16            loss = criterion(net(X_batch), y_batch)
17            loss.backward()
18            optimizer.step()
Code 7: Deep Learning with Differential Privacy
Additional examples are available in our documentation (https://koukyosyumei.github.io/AIJack/).
Acknowledgement
We acknowledge every contributor for their support of this project. The comprehensive list of contributors can be found in our GitHub repository.